# Neuromorphic detection and cooling of microparticles in arrays

### v3

@@ Methods
Figure 5 shows the experimental setup surrounding our linear Paul trap. The trap consists of 4 parallel cylindrical trapping electrodes forming a square, with two coaxial cylindrical endcap electrodes, one of which we call the control electrode which is used for feedback control. The distance from trap centre to the surface of the -diameter trapping electrodes is , and one opposing-pair are driven with 360 V amplitude at 1 kHz. Of the other pair of the four electrodes, the lower one has a constant voltage of 3V applied to minimize single-particle micromotion. The two endcaps are diameter and separated by , with a slight misalignment along the axis which causes particles to be trapped along a diagonal in the plane. The proximity of the electrodes to the centre of the trap means that a voltage applied to either one will create a field with a significant component in all three axes, enabling 3D control with a single electrode.
Laser induced acoustic desorption (LIAD) [nikkhou2021] is used to launch microparticles into the trap at a pressure of mbar and then the pressure is lowered to carry out the experiments presented in this manuscript. We typically trap particles of positive charge, ranging from to .
To image the particles, 18 mW of laser light of wavelength nm is weakly focused onto the array. The scattered light is imaged onto an event-based camera (Prophesee EVK1 -Gen3.1 VGA (camera sensor: Prophesee PPS3MVCD, pixels)) using a long working distance microscope. The camera is precisely calibrated using a method outlined in detail in [ren2022event]. When dealing with multiple particles, calibration is performed via displacing a translation stage on which our imaging system is mounted by a known amount.
@@
The data pipeline is shown in Fig. 6. The operation of the EBC is described in detail in [ren2022event]. The generic tracking algorithm (GTA) of the EBC outputs 2D position data for each object it detects. The EBC is communicated with via a Python script, which separates this 2D information into two 1D data streams for each object. The script passes each data stream to one of three FPGA systems (Red Pitaya STEMlab 125-14) to output the position of each particle. The FPGA clock is synchronized to the clock of the EBC to ensuring timing consistency. Code on the FPGA computes the velocity from the position data, and adds a variable gain and phase to the signal to generate the feedback signal. There is a latency of 10 ms in this pipeline, see Supplementary Materials S4. EBCs are available with FPGA systems on the camera hardware, which will significantly reduce this latency.
@@ Acknowledgments
In this Supplemental Material we provide technical details on the experiment and data analysis, including: S1 Coordinate systems; S2 Mode identification for multiple particles; S3 Extended information for the four-particle dataset; S4 Latency in the feedback loop; S5 Power consumption of the event based camera.
@@ Appendix 1 Coordinate systems
Our setup has two coordinate systems: the Paul trap coordinates , and the camera coordinates , as shown in Supplementary Figure 1 and Fig. 1 of the manuscript. The Paul trap coordinates define the oscillation axes of the levitated particles. The -axis is defined along the diagonal between the two trapping electrodes held at a DC voltage (see Methods in the main manuscript), the -axis along the diagonal between the two trapping electrodes with an AC voltage, and the -axis along the axis parallel to the endcap electrodes. In the camera frame, the trap - and -axes are projected onto the camera’s -axis, and the camera’s -axis is parallel to the -axis. Therefore, we can capture the 3D motions of the levitated particles with the camera’s 2D image.
@@ Appendix 2 Mode identification for multiple particles
To identify collective modes due to interactions between the charged particles, we calculate the cross-spectral density (CSD), which picks-out only the correlated spectral components [penny2000signal]. As shown in Supplementary Figure 3(a), none of the six modes exhibit significant coupling, since the separation between the particles (m) is too large.
@@ Appendix 3 Extended information for the four-particle dataset
Here we investigate the collective modes of the four levitated microparticles case shown in Fig. 2 of the manuscript. As discussed in the Supplementary Section S2, the collective modes can be found via a CSD analysis. We show CSDs for all pairs of particles in Supplementary Figure 4, and conclude that the four levitated particles are coupled along the Paul trap -axis (14 Hz, 25 Hz, 32 Hz and 35 Hz) due to their layout in the trap [cf. Supplementary Figure 1]. It is expected that four coupled oscillators have four coupled modes, but the nature of them is intricate and cannot be simply attributed as common or breathing modes unlike the two particle case [penny2023sympathetic, Watson_2020].
@@
We report the signal-to-noise ratios (SNRs) of our detection for each of the four particles in Supplementary Table 1. We can see the strongest SNR is reported for particle P4, and it gradually reduces for particles further away from this. This is because we use a single laser for illumination focusing on P4. In single-particle experiments, we achieve a SNR up to dB [ren2022event]. The SNR can be improved in future experiments using a higher-power illumination system which can be expanded to uniformly illumnate many particles. Alternatively, separate beams could be used for each particle, via e.g. a spatial light modulator [obata2010multi, rieser2022tunable].
@@ Appendix 4 Latency in the feedback loop
The delay of applied control signal has two effects (following Ref. [debiossac2020]). Firstly, it deterministically shifts the phase between the mechanical system and the feedback signal, leading to oscillatory behaviour in the correlations and thus a transition between heating and cooling. Secondly, it causes stochastic dephasing of the mechanical motion relative to the feedback signal as the delay becomes large, which reduces the strength of the correlations.
From the experimental results of Ref. [debiossac2020], correlation functions of the microparticle centre-of-mass motion as a function of delay remain high for short delays (on the order of a few oscillation periods) when the dynamics are underdamped (as in our system). Since our levitated oscillators have oscillation frequencies below 100 Hz, the latency of 10 ms corresponds to a single period of delay, and has little consequence for our cooling protocol.

### v4

@@ Methods
Figure 5 shows the experimental setup surrounding our linear Paul trap. The trap consists of four parallel cylindrical trapping electrodes forming a square, with two coaxial cylindrical endcap electrodes, one of which we call the control electrode which is used for feedback control. The distance from trap centre to the surface of the -diameter trapping electrodes is , and one opposing-pair are driven with 360 V amplitude at 1 kHz. Of the other pair of the four electrodes, the lower one has a constant voltage of 3V applied to minimize single-particle micromotion. The two endcaps are diameter and separated by , with a slight misalignment along the axis which causes particles to be trapped along a diagonal in the plane. The proximity of the electrodes to the centre of the trap means that a voltage applied to either one will create a field with a significant component in all three axes, enabling 3D control with a single electrode.
Laser induced acoustic desorption (LIAD) [85] is used to launch microparticles into the trap at a pressure of mbar and then the pressure is lowered to carry out the experiments presented in this manuscript. We typically trap particles of positive charge, ranging from to .
To image the particles, 18 mW of laser light of wavelength nm is weakly focused onto the array. The scattered light is imaged onto an event-based camera (Prophesee EVK1 -Gen3.1 VGA (camera sensor: Prophesee PPS3MVCD, pixels)) using a long working distance microscope. The camera is precisely calibrated using a method outlined in detail in [89]. When dealing with multiple particles, calibration is performed via displacing a translation stage on which our imaging system is mounted by a known amount.
@@
The data pipeline is shown in Fig. 6. The operation of the EBC is described in detail in [89]. The generic tracking algorithm (GTA) of the EBC outputs 2D position data for each object it detects. The EBC is communicated with via a Python script, which separates this 2D information into two 1D data streams for each object. The script passes each data stream to one of three FPGA systems (Red Pitaya STEMlab 125-14) to output the position of each particle. The FPGA clock is synchronized to the clock of the EBC to ensuring timing consistency. Code on the FPGA computes the velocity from the position data, and adds a variable gain and phase to the signal to generate the feedback signal. There is a latency of 10 ms in this pipeline, see Supplementary Materials S4. EBCs are available with FPGA systems on the camera hardware, which will significantly reduce this latency.
@@ Acknowledgments
In this Supplemental Material we provide technical details on the experiment and data analysis, including: S1 Coordinate systems; S2 Mode identification for multiple particles; S3 Extended information for the four-particle dataset; S4 Latency in the feedback loop; S5 Power consumption of the event based camera; S6 Bath temperatures for all particles; S7 Cooling limit and the noise squashing; S8 Scalability of multi-mode cooling.
@@ Appendix 1 Coordinate systems
Our setup has two coordinate systems: the Paul trap coordinates , and the camera coordinates , as shown in Supplementary Figure 1 and Fig. 1 of the manuscript. The Paul trap coordinates define the oscillation axes of the levitated particles. The -axis is defined along the diagonal between the two trapping electrodes held at a DC voltage (see Methods in the main manuscript), the -axis along the diagonal between the two trapping electrodes with an AC voltage, and the -axis along the axis parallel to the endcap electrodes. In the camera frame, the trap - and -axes are projected onto the camera’s -axis, and the camera’s -axis is parallel to the -axis. Therefore, we can capture the 3D motion of the levitated particles with the camera’s 2D image.
@@ Appendix 2 Mode identification for multiple particles
To identify collective modes due to interactions between the charged particles, we calculate the cross-spectral density (CSD), which picks-out only the correlated spectral components [86]. As shown in Supplementary Figure 3(a), none of the six modes exhibit significant coupling, since the separation between the particles (m) is too large.
@@ Appendix 3 Extended information for the four-particle dataset
Here we investigate the collective modes of the four levitated microparticles case shown in Fig. 2 of the manuscript. As discussed in the Supplementary Section S2, the collective modes can be found via a CSD analysis. We show CSDs for all pairs of particles in Supplementary Figure 4, and conclude that the four levitated particles are coupled along the Paul trap -axis (14 Hz, 25 Hz, 32 Hz and 35 Hz) due to their layout in the trap [cf. Supplementary Figure 1]. It is expected that four coupled oscillators have four coupled modes, but the nature of them is intricate and cannot be simply attributed as common or breathing modes unlike the two particle case [87, 88].
@@
We report the signal-to-noise ratios (SNRs) of our detection for each of the four particles in Supplementary Table 1. We can see the strongest SNR is reported for particle P4, and it gradually reduces for particles further away from this. This is because we use a single laser for illumination focusing on P4. In single-particle experiments, we achieve a SNR up to dB [89]. The SNR can be improved in future experiments using a higher-power illumination system which can be expanded to uniformly illuminate many particles. Alternatively, separate beams could be used for each particle, via e.g. a spatial light modulator [90, 91]. The charge to mass ratios of the four levitated particles are provided in Supplementary Table 2, from which we can observe that the four charges are different, since the masses are the same to within 10% (manufacturer specified).
@@ Appendix 4 Latency in the feedback loop
The delay of applied control signal has two effects (following Ref. [92]). Firstly, it deterministically shifts the phase between the mechanical system and the feedback signal, leading to oscillatory behaviour in the correlations and thus a transition between heating and cooling. Secondly, it causes stochastic dephasing of the mechanical motion relative to the feedback signal as the delay becomes large, which reduces the strength of the correlations.
From the experimental results of Ref. [92], correlation functions of the microparticle centre-of-mass motion as a function of delay remain high for short delays (on the order of a few oscillation periods) when the dynamics are underdamped (as in our system). Since our levitated oscillators have oscillation frequencies below 100 Hz, the latency of 10 ms corresponds to a single period of delay, and has little consequence for our cooling protocol.
@@ Appendix 5 Power consumption of the event based camera
Appendix 6 Bath temperatures for all particles
Due to voltage noise from the amplifiers driving our Paul trap, the equilibrium temperature of our particles without cooling ranges from approximately K. This temperature varies depending on their charge and spatial location within the trap. Here we model the electrical noise as a white noise bath and assume the particles are trapped in a quadratic potential. Under the assumption that the equipartition theorem holds, the different bath temperatures used in the main text figures are listed in Supplementary Table 3. In Fig. 4(c), P2 is located closest to the trap center and therefore experiences the least electric field noise, resulting in the lowest bath temperature. In contrast, P1 and P3 are farther from the trap center, and their bath temperatures are approximately an order of magnitude higher than that of P2. A similar trend can be observed in Fig. 4(b), where P1, being farther from the trap center, also exhibits a higher temperature than P2.
| Fig3(b) | Fig3(c) | Fig3(d) | Fig4(a) | Fig4(b) | Fig4(c) | |
P1:
P2: |
P1:
P2: P3: |
Appendix 7 Cooling limit and the noise squashing
At a certain pressure (which determines the gas damping rate ), to achieve maximal cooling the optimal feedback rate can be obtained from equation (2) in the main manuscript by taking [93], which gives
with the optimal feedback rate
Note that is -independent.
The relation between the feedback rate and the cooling temperature is shown in Supplementary Figure 6(a), where we plot equation (2) in the main manuscript at three different pressures. The dots in the plot are the optimal feedback rates and the optimal temperatures corresponding to the respective pressure, computed from Supplementary Equations (1) and (2). As we can see, the optimal feedback rate gets smaller as the pressure goes down.
At a certain pressure, a larger feedback rate results in a lower system temperature (hence a better cooling), until it pass the the optimal value in which case the feedback signal instead heats up the system. This is because the increasing amplitude of the feedback signal also amplifies the noise fed back to the particle, preventing further cooling. The corresponding PSD also appears inverted below the noise floor, a phenomenon known as the noise squashing [94]. In Supplementary Figure 6(b) we demonstrate this by setting the feedback rate Hz and reducing the pressure. Correspondingly, the PSDs of the measured motion changes from unsquashed (blue line, ), to mixed (orange line, ), and finally to squashed (green line, ). Furthermore, the squashed shape is asymmetrical due to the delay in the signal [95]. This makes a small change to the fitting PSD function, as shown in the following:
Suppose we consider a time delay in the feedback signal , the equation of motion reads
Take the Fourier transform with and , we get the particle’s detected motion in the frequency domain
We can calculate the PSD in this situation by taking and assume that two noises and are uncorrelated, which gives the PSD of the measured position in the in-loop detector:
This signal delay can cause a small asymmetry that is visible in the squashed PSD from in-loop detection. In this case, the experimental PSD of particle’s measured motion can be correctly fitted by Supplementary Equations (3), as shown in Supplementary Figure 6(b).
Appendix 8 Scalability of multi-mode cooling
Here we discuss the fundamental limit on the number of modes we can cool with the technique presented in our manuscript. We assume is bounded by the bandwidth of our detection Hz and the minimum resolvable linewidth , such that . Then, to estimate the number of modes, we need to find the minimum linewidth .
The contribution to this limit from the experimental setup comes mainly from filtering. Considering the time duration for signal demodulation, there is a bandwidth limit for filters to pick out certain frequency signal which will set a limit to . A PLL has a very narrow filter bandwidth, for example a Zurich Instruments HF2LI has a minimum bandwidth of Hz and can filter 6 peaks at once. Therefore, with PLL it is possible to pick up feedback signal at mHz resolution.
On the other hand, since we use cold damping to cool the particles, the particles inevitably experience an additional damping rate, which broadens the linewidth and makes the modes harder to distinguish. According to Eq. (2) in the manuscript, by solving , we can obtain an optimal at a given pressure (See Supplementary Equations (2)). This optimal would be the dominant factor in determining the scaling number, whose value reduces as pressure goes down.
If we take the above argument as a rule of thumb, we can estimate the number of modes as follows. Suppose we start cooling at mbar in experiment, at which the particle’s linewidth is about Hz, the optimal feedback rate is Hz. This limits the capacity of our setup to be modes, which can be improved at high vacuum.